as:

A charity bingo game costs $2 per round and has a $13 entry fee for an adult. It costs $3 per round and a $10 entry fee for a child. For how many rounds of bingo are the costs the same for an adult and a child?

AndrewArni:

What are the answer choice?

AndrewArni:

Yep

jhonyy9:

@as you give the right answer but please explain why are 3 rounds the correct answer to understand everybody

darkknight:

So to solve this we can use a systems of equations. The entry fee for an adult is $13, which we can use as our y intercept in the equation a(x). Our slope will be $2 because that is how much it costs per round. So our equation for the adults is a(x) = 2x+13 For our entry fee for a child is $10, (our y-intercept). And our slope is $3. So c(x) = 3x+10. To find when the rounds cost the same. We set them equal to each other and then solve a(x) = c(x) 2x+13 = 3x+10 2x - 2x +13 = 3x - 2x+10 13 = x+10 x = 3